Twins Joanie and Tony each start out with the same number of cents.
Joanie buys 1 candy bar and has 70 cents left.
Tony buys 3 candy bars at the same price and has 20 cents left
How many cents did Joanie start with?

Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

x equals the price of a candy bar.
y equals the amount of money they both started out with.

your equations are:

for joanie, y - x = 70

for tony, y - 3x = 20

these are two equations that need to be solved simultaneously.

subtract the second equation from the first to get:

2x = 50

solve for x to get x = 25

for joanie, y - 25 = 70.

solve for y to get y = 95 cents.

for tony, y - 3 * 25 = 20 which becomes y - 75 = 20.

solve for y to get y = 95 cents.

they both started out with 95 cents.

the price of one candy bar was 25 cents.

95 - 25 = 70 which is what joanie had left.

95 - 3 * 25 = 20 which is what tony had left.





Answer by greenestamps(13200)



The additional cost of the extra 2 candy bars was (70-20) = 50 cents. So each candy bar costs 50/2 = 25 cents.

Since Joanie bought 1 candy bar and had 70 cents left, the amount she started with was 70+25 = 95 cents.